Could you help me find what x equals in these equations? I appear to have forgotten how to solve for x...
4x+3=0 and x-2=0

Answer by Cromlix(4381)

Hi there,
4x + 3 = 0
4x = -3
x = -3/4
.....
x - 2 = 0
x = 2
........
Hope this helps :-)